A boat travels 45 miles downstream in the same time it takes to travel 30 miles upstream. If the speed of the boat in still water is 12 mph what's the rate of the current?
Answer by mananth(16949)

Boat speed = 12 mph
current speed = x mph
wind speed
speed against current 12-x
speed with current 12+x
Distance against current 30 mph
Distance with current 45 mph
t=d/r time upstream = time downstream
30/(12-x)=45/(12+x)
30(12+x)=45(12-x)
360+30x =540-45 x
45x+30x =180
75x=180
/75
x=2.4 mph current speed